The vapour pressures of pure liquids . A and B at 300 K are 4.5×10⁴ Pa and 6.5× 10⁴ Pa respectively. calculate the mole fractions of of A and B in vapour and liquid phases of a solution when the liquid equilibrium total vapour pressure of the binary liquid solution is 5.5× 10⁴ Pa at 300 K . Assume that the liquid and the vapour are ideal .
1
Expert's answer
2019-10-31T07:29:12-0400

Let mole fraction of A in liquid phase is "x_a" and B in liquid phase is "x_b=1-x_a"

Then according to question,


"0.55=0.45x_a+0.65(1-x_a)"

"0.2x_a=0.1\\or\\x_a=0.5"

"x_b=1-x_a=0.5"

Composition in vapour phase will be-

"p_a=0.45x_a=0.225\ bar\\p_b=0.65x_b=0.325\ bar"

Mole fraction of A in vapour phase , "y_a=\frac{0.225}{0.55}=0.4090"

Mole fraction of B in vapour phase,"y_b=\frac{0.325}{0.550}=0.5909"